OpenStudy (anonymous):

A coin is tossed 2500 times. There is about a 68% chance that the percent of heads is in the range 50% plus or minus 0.5% 1% 1.5% 2% 2.5%

OpenStudy (dumbcow):

im thinking its 1% if this is binomial distribution... \[\sigma = \sqrt{np(1-p)} = \sqrt{2500*.5*.5} = 25\] 25 is 1% of 2500

OpenStudy (anonymous):

yes ,it is right ans